TAX in Canada

Paige Turner is a sales manager employed in Ontario. She used her car and her cell phone for employment related purposes during 2019. She claimed the following expenses on her T777, Stmt of Employment Expenses, for 2019 (all HST taxable amounts include applicable HST):

•Cell phone charges  $1,391

•Gas, maintenance and car repairs $10,886

•Insurance on car $539

The car was purchased during 2019. She did not own a car prior to this purchase.

Calculate the HST/GST rebate.
Answer #1

Hst rate in 2019 is 13%

Then

Total value = 1391 +10886 = 12277

* she did not own a car then insurance on car is not taken. If CCA value is given it is included.

Hst/gst rebate = value *rate/100+rate

= 12277*13/113 = 1412.39

*If CCA valus is it added to total value then compute hst/gst rebate as same formula.


answered by: Zahidul Hossain